# Parameter Efficient Continual Learning with Dynamic Low-Rank Adaptation

### v2

@@ 1 Introduction
Learning forms the foundation for intelligent systems to adapt to dynamic environments in response to external changes. Unlike conventional deep neural networks (DNNs) that are designed for static data distribution, Continual Learning (CL) is characterized by learning from sequential, dynamic data distribution [67]. However, CL has remained a significant challenge for DNNs due to the issue of catastrophic forgetting: A phenomenon in which adaptation to new tasks results in significant deterioration of consolidated knowledge [51, 14, 15, 55, 6]. A number of approaches such as rehearsal-based (e.g., [10, 5]), weight-regularization based (e.g., [37]) and parameter-isolation based (e.g., [57]) have been proposed in the literature to address the problem of catastrophic forgetting in CL. Although CL is primarily intended to ensure the resource efficiency of model updates [67], the reduction in catastrophic forgetting in these approaches is reliant on increases in computational and memory footprint due to model surrogates and memory buffer. Extrapolating forward this reliance in real world scenarios is both economically and environmentally infeasible.
@@
The increased availability of pre-trained models that demonstrate strong generalization capabilities has led researchers to discover early exploratory methods [17, 34, 22, 73], facilitating more efficient integration of new knowledge in CL systems. However, continuous fine-tuning on multiple downstream tasks compromises pre-trained model’s generalization capabilities [46]. Recently, Parameter Efficient Fine-Tuning (PEFT) CL techniques [45, 78] are on the rise as they provide lightweight training schedule while avoiding degradation of consolidated knowledge in pre-trained models. However, it is not clear how one can select the size of their rank (in case of LoRA [29]) while their performance is very sensitive to rank selection [62]. Although choosing a higher rank usually results in optimal performance, it significantly increases the number of learnable parameters in longer task sequences when using large pre-trained models. Task arithmetic [33], which enables model compression using simple arithmetic operations with task vectors (difference between pre-trained and fine-tuned weights) [13], allows for selection of higher rank during training without any overhead during inference. Task arithmetic assumes that fine-tuning a pre-trained model pushes the weights away towards a semantically relevant direction in the weight space, thus enabling model compression using simple arithmetic operations with task vectors [13]. However, task arithmetic suffers from number of shortcomings including finding optimal scaling coefficients and suboptimal performance in the absence of a memory buffer [79].
@@
Typically, when using arbitrary weights, a higher rank usually facilitates learning new tasks but tends to increase forgetting with significant increase in number of parameters, while lower-rank mitigates forgetting with limited adaptation [48, 8]. There often exists a balance between plasticity and stability at a moderately small (but not too small) rank, which may enhance the benefits of CL [48]. To this end, we introduce PEARL (Parameter-Efficient Adaptive Ranking for lifelong learning with LoRA), a rehearsal-free, parameter-isolation-based PEFT vision CL framework that mitigates catastrophic forgetting through dynamic low-rank adaptation. Specifically, PEARL leverages the information contained in the reference task weights  [^1] We use reference task interchangeably with pre-training when available. Otherwise, it refers to the first task in CL. and dynamically determines the rank of task-specific, layer-specific LoRA components during training based on the current task’s proximity to reference task weights in parameter space. With our dynamic threshold criterion, a close proximity would result in a lower rank allocation and maximum information re-use while a distant proximity results in a larger rank selection, thereby finding an optimal balance between learning and forgetting. To demonstrate the versatility of PEARL, we perform extensive analyses across various datasets, multiple CL scenarios and different architectures, an overview of which can be found in Appendix A.1. Our results demonstrate that PEARL variants consistently outperform the evaluated baselines in all scenarios by a huge margin. Our contributions are as follows:
@@ 2 Related Works
Exemplar and Non-Exemplar Based CL: Experience Rehearsal (ER), which stores and replays samples from previous tasks, has remained a leading approach to addressing the issue of catastrophic forgetting in CL. Several works build upon ER: DER++ [10] proposed a consistency regularization technique to enhance stability and generalization. Guided by the GWT, TAMiL [7] and AGILE [3] incorporate information reuse and task-specific, parameter-efficient adapters that re-project the backbone features onto a global workspace. However, these approaches face additional computational and memory overhead due to model surrogates and experience rehearsal. Since storing samples from previous tasks is not always feasible, non-exemplar-based approaches are gaining traction. Non-Exemplar Class-IL (NECIL) is particularly important in scenarios where data confidentiality is crucial due to privacy or security concerns, and where data storage has a limited lifespan. Early contributions to this domain include LwF [44] and EWC [37], both of which employ regularization techniques to reduce catastrophic forgetting. More recent methods, such as PASS [81] and IL2A [80], have further advanced NECIL by generating prototypes for previous classes without the need to retain the original images. SSRE [82] introduced a reparameterization method to balance old and new knowledge, along with additional self-training that utilizes external data as an alternative to exemplars. FeTrIL [53] proposed a framework that combines a fixed feature extractor with a pseudo-feature generator using geometric transformations to achieve a better stability-plasticity balance. FeCAM [24] addressed the limitations of Euclidean distance when feature distributions are non-stationary and suggested using anisotropic Mahalanobis distance as a substitute for Euclidean distance when employing a frozen feature extractor for prototype classification. Overall, the NECIL benchmark presents a significant challenge as it prohibits the storage of any samples from previous tasks while requiring effective performance in a Class-IL setting. PEARL also aligns with the NECIL benchmark as it operates without any experience rehearsal.
@@
Parameter Efficient Fine-Tuning (PEFT) techniques that have emerged in the space of Large Language Models have been quite pivotal in optimizing model performance with little memory and computational overhead. Seminal works such as Adapter modules [28], Prompt Tuning [9], BitFit [74], LoRA [29], and DoRA [47] have demonstrated the efficacy of selective fine-tuning all the while preserving model generalization and enhancing model adaptability. Specifically, LoRA introduces two compact, trainable low rank matrices into each layer’s weight matrix, allowing model to adapt to new tasks with minimal memory and computational footprint. Within CL, several approaches leverage PEFT techniques to mitigate catastrophic forgetting: Prompt-based methods such as L2P [72] and DualPrompt [70] introduce task-specific prompts to facilitate learning new tasks while preserving consolidated knowledge. S-Prompt [68] and CODA-Prompt [60] extend these works by employing structural prompts to map discriminative relationships and through Schmidt orthogonalization respectively. In order to maintain an optimal balance between computational complexity and accuracy, adapter-based methods such as SEMA [65] and DIA [41] propose self-expanding, dynamic integration of task-specific adapters. Similarly, EASE [78] creates task-specific sub-spaces by employing lightweight adapter modules for each new task, thereby maintaining resource efficiency. In contrast to adapter-based methods, InfLoRA [45] inserts a small number of LoRA parameters to re-parameterize the pre-trained weights and shows that fine-tuning these inserted parameters is equivalent to fine-tuning the pre-trained weights within a subspace. C-LoRA [59] incorporates a self-regularization strategy, which penalizes updates to parameters already modified by previous tasks. This mechanism helps preserve prior knowledge and enhances stability across sequential tasks. While these approaches have shown relative success, their implementation is tailored specifically to pre-trained vision transformers and do not consider task dynamics, limiting the potential for information reuse and dynamic resource allocation.
@@
Task Arithmetic [33], centered around task vectors, offers a new paradigm for steering the behavior of pre-trained / reference network towards multiple downstream tasks using task similarity. Inspired by recent works on weight interpolation [19, 32, 1], task vectors are defined to be the element-wise difference between pre-trained / reference network weights before and after fine-tuning on a particular task. A task vector specifies a direction in the weight space of a pre-trained model, such that movement in that direction improves performance on a given task [33]. As these task vectors can be modified and combined together in simple arithmetic operations, new tasks can be learnt via addition while interfering tasks can be forgotten via negation. Inspired by these findings, Chitale et al. [13] proposed an approach to continually train transformer-based vision models using low-rank adaptation and task arithmetic. Although task arithmetic is quite lucrative for its simplicity and intuitiveness, it suffers from several shortcomings such as finding optimal scaling coefficients and sub-optimal performance in the absence of memory buffer to finetune the network after weight aggregation [79].
@@ 3.1 Information aggregation during reference task training
where ${b}$ is a training batch, $\mathcal{L}_{ce}$ is cross-entropy loss, $\mathcal{D}_{r}$ is the data distribution for the reference task, and $\sigma(.)$ is the softmax function. Since pre-trained models demonstrate strong generalization capabilities, we expect to aggregate generic information that can be re-used across tasks later during CL. In scenarios where pre-training on a large corpus of data is not a possibility, the first task in a CL setup can also be used as a reference task.
@@ 3.2 Information re-use and dynamic resource allocation during CL
[…]
@@
where $\mathcal{T}$ represents a normalized squared error between current task and reference task weights, and $\circ$ represents element-wise multiplication. Through $\mathcal{T}$, we approximate the proximity of the current task with respect to the reference task in the parameter space. Once $\mathcal{T}$ is computed, the dynamic rank for the task-specific, layer-specific LoRA can be found as follows:
$k=\arg\max_{j}\left(\frac{\sum_{i=1}^{j}\mathbf{\sigma}_{i}^{2}}{\sum_{i=1}^{r}\mathbf{\sigma}_{i}^{2}}\geq\mathcal{T}_{\mu}\right)+1$ (4)
where $\mathcal{T}_{\mu}$ represents the scalar mean of the $\mathcal{T}$. Equation 4 ensures that the k is bounded and selects the rank in proportion to the cumulative explained variance of singular values. Subsequently, the low-rank matrices can be initialized from the SV-Decomposition of $\mathcal{W}^{t}_{c}$ (flattened and re-shaped to 2D if necessary) as follows:
@@
Proposition 3.2 (Boundedness of $\mathcal{T}$)
Let $\mathcal{W}^{r}$ be a non-empty reference task matrix such that $\exists\;i,j$ for which $\mathcal{W}^{r}_{ij}\neq 0$. Then, the dynamic threshold $\mathcal{T}$ is bounded between 0 and 1.
Proof: By the properties of real numbers, it follows that $\mathcal{T}\geq 0$ as element-wise multiplication and consequent sum of a matrix is always non-negative. Since $\mathcal{W}^{t}_{c}=\mathcal{W}^{t}-\mathcal{W}^{r}$, we can re-write Equation 3 denominator as $\sum(\mathcal{W}^{t}_{c}\circ\mathcal{W}^{t}_{c})+2\sum(\mathcal{W}^{r}\circ\mathcal{W}^{t}_{c})+2\sum(\mathcal{W}^{r}\circ\mathcal{W}^{r})$ Since the all the terms in the denominator are also non-negative, it trivially follows that denominator is greater than or equal to numerator resulting in $\mathcal{T}\leq 1$. Therefore, the dynamic threshold $\mathcal{T}$ will be bounded between 0 and 1 when reference tasks weights are not all zeroes.
@@ 3.3 Putting it all together
Table 1: Comparison of Class-IL performance of exemplar-based and NECIL approaches across different datasets and number of tasks. Baselines use randomly initialized ResNet-18 as a backbone. Exemplar-based approaches use a buffer size of 200. NECIL results are taken from [25], while for the rest, we report an average of 3 runs. Best results are marked in bold while second best are underlined.
@@
DER++ [10] NeurIPS 2020 28.8 47.89 25.42 43.94 17.75 35.89 10.15 27.26
TAMiL [7] ICLR 2023 41.43 58.09 32.20 49.87 26.32 43.91 20.46 40.94
AGILE [3] CoLLAs 2024 45.73 57.66 33.23 42.88 20.14 43.79 20.19 43.81
LwF [44] ECCV 2016 45.35 61.94 26.14 46.14 38.81 49.70 27.42 38.77
IL2A [80] NeurIPS 2021 - - 31.70 48.40 - - 25.30 42.01
PASS [81] CVPR 2021 49.75 63.39 37.78 52.18 36.44 48.64 26.58 38.65
SSRE [82] CVPR 2022 42.39 56.57 29.44 44.38 30.13 43.20 22.48 34.93
FeTrIL [53] WACV 2023 45.11 60.42 36.69 52.11 29.91 43.99 23.88 36.35
FeCAM [24] NeurIPS 2024 47.28 61.37 33.82 48.58 25.62 39.85 23.21 35.32
@@ 4 Results
Experimental settings: We evaluate PEARL across two distinct scenarios of CL: Class-IL and Task-IL, a complete overview of which can be found in Appendix A.1. To facilitate this evaluation, we consider a comprehensive array of baseline methods: (i) exemplar-based approaches that involve information reuse and task-specific adapters (e.g., TAMiL, AGILE), (ii) NECIL approaches (e.g., PASS, FeTriL, FeCAM), (iii) parameter-isolation approaches that operate within a fixed capacity (e.g., NISPA) and those that expand beyond it (e.g., PNN, PAE), and (iv) PEFT approaches for CL (e.g., L2P, C-LoRA, InfLoRA, SEMA, ADA). We also provide a justification for the choice of these baseline in Appendix A.2. The first three categories utilize randomly initialized ResNet-18 as their backbone, whereas PEFT approaches leverage the ViT-Base-16 model pre-trained on ImageNet-21k. It is important to note that pre-training is often a luxury for smaller, custom models. Consequently, in accordance with the training protocols of the baseline methods, we employ randomly initialized convolutional models, while the ViT model is pre-trained. To demonstrate the versatility of PEARL across different architectures, we explore three distinct backbones: (i) Convolutional networks such as ResNet-18 (R18) and a ResNet-18-like architecture incorporating Blueprint Separable Convolutions (BSC) [43], and (ii) the Vision Transformer (ViT) Vit-Base-16 pre-trained on ImageNet-21k. For each subsequent task, we adapt convolutional layer, point-wise convolutional layer and linear layer projecting key in self-attention using dynamic low-rank adaptation in each of these backbones respectively. More information on backbones and their settings can be found in Appendix C.
@@
Comparison with convolutional CL approaches: Table 1 presents a comparison of PEARL against exemplar-based and NECIL approaches across four different Class-IL settings. We report results for both of our convolutional variants. Since the backbone is randomly initialized without pre-training, PEARL convolutional variants use the first CL task as the reference task and assign LoRA components for subsequent tasks based on their proximity to the first task in the parameter space (Algorithm 2). As shown, PEARL (R18) outperforms all considered baselines without relying on experience rehearsal while exhibiting modest model growth with an increasing number of tasks (see Table 11). Although a larger buffer size might enhance the performance of rehearsal-based approaches, any improvements would likely be outweighed by the increased computational and memory costs [64]. Additionally, NECIL approaches do not always re-use information across tasks, which often leads to subpar performance. In contrast, PEARL capitalizes on the information contained in the first task for every subsequent task, achieving an optimal balance between performance and model size. While the absence of a pre-trained model may limit the generalizability of the first task, the framework’s ability to reuse information and dynamically allocate resources allows PEARL to efficiently assimilate information from both the first and current tasks, resulting in superior performance across tasks.
@@
Table 2: Class-IL performance comparison on ImageNet-R across 5, 10, and 20 task sequences. Baseline results with standard deviation are taken from [45] and that of PEARL are averaged over 3 runs with errors indicating standard deviation across these runs. The rest of the results are taken from their respective manuscripts. Best results are highlighted in bold. SeqLoRA is a sequential learning baseline which fine-tunes LoRA components on multiple tasks without any additional support.
@@
L2P [72] CVPR 2022 64.13 $\pm$ 0.78 68.66 $\pm$ 0.41 62.54 $\pm$ 0.24 67.98 $\pm$ 0.27 57.92 $\pm$ 0.28 64.57 $\pm$ 0.29
DualPrompt [71] ECCV 2022 67.88 $\pm$ 0.17 71.16 $\pm$ 0.31 65.41 $\pm$ 0.52 69.39 $\pm$ 0.43 61.00 $\pm$ 0.72 65.80 $\pm$ 0.67
CODA-P [60] CVPR 2023 73.09 $\pm$ 0.21 76.91 $\pm$ 0.21 71.47 $\pm$ 0.35 75.82 $\pm$ 0.29 67.28 $\pm$ 0.30 72.34 $\pm$ 0.17
C-LoRA [59] CoRR 2023 75.85 $\pm$ 0.31 78.85 $\pm$ 0.34 71.89 $\pm$ 0.45 75.33 $\pm$ 0.28 65.71 $\pm$ 0.60 70.63 $\pm$ 0.85
LAE [21] ICCV 2023 73.84 $\pm$ 0.14 77.29 $\pm$ 0.45 71.70 $\pm$ 0.39 76.71 $\pm$ 0.10 66.98 $\pm$ 0.35 73.72 $\pm$ 0.05
InfLoRA [45] CVPR 2024 77.52 $\pm$ 0.37 82.01 $\pm$ 0.12 75.65 $\pm$ 0.14 80.82 $\pm$ 0.24 71.01 $\pm$ 0.45 77.28 $\pm$ 0.45
EASE [78] CVPR 2024 70.80     - 73.37     - 78.80     -
SEMA [65] CVPR 2025 79.78 84.75 78.00 83.56 74.53 81.75
DIA [41] CVPR 2025     -     - 79.03 85.61 76.32 83.51
TSVD [52] ICLR 2025 73.58     - 74.55     - 79.13     -

### v3

@@ 1 Introduction
Learning forms the foundation for intelligent systems to adapt to dynamic environments in response to external changes. Unlike conventional deep neural networks (DNNs) that are designed for static data distribution, Continual Learning (CL) is characterized by learning from sequential, dynamic data distribution [wang2023comprehensive]. However, CL has remained a significant challenge for DNNs due to the issue of catastrophic forgetting: A phenomenon in which adaptation to new tasks results in significant deterioration of consolidated knowledge [masana2022class, de2021continual, ditzler2015learning, robins1995, pmlr-v199-bhat22a]. A number of approaches such as rehearsal-based (e.g., [buzzega2020dark, bhat2022consistency]), weight-regularization based (e.g., [kirkpatrick2017overcoming]) and parameter-isolation based (e.g., [rusu2016progressive]) have been proposed in the literature to address the problem of catastrophic forgetting in CL. Although CL is primarily intended to ensure the resource efficiency of model updates [wang2023comprehensive], the reduction in catastrophic forgetting in these approaches is reliant on increases in computational and memory footprint due to model surrogates and memory buffer. Extrapolating forward this reliance in real world scenarios is both economically and environmentally infeasible.
@@
The increased availability of pre-trained models that demonstrate strong generalization capabilities has led researchers to discover early exploratory methods [douillard2022dytox, jeeveswaran2023birt, gao2024consistent, yang2024generating], facilitating more efficient integration of new knowledge in CL systems. However, continuous fine-tuning on multiple downstream tasks compromises pre-trained model’s generalization capabilities [liu2024parameter]. Recently, Parameter Efficient Fine-Tuning (PEFT) CL techniques [liang2024inflora, zhou2024expandable] are on the rise as they provide lightweight training schedule while avoiding degradation of consolidated knowledge in pre-trained models. However, it is not clear how one can select the size of their rank (in case of LoRA [hu2021lora]) while their performance is very sensitive to rank selection [valipour2022dylora]. Although choosing a higher rank usually results in optimal performance, it significantly increases the number of learnable parameters in longer task sequences when using large pre-trained models. Task arithmetic [ilharcoediting], which enables model compression using simple arithmetic operations with task vectors (difference between pre-trained and fine-tuned weights) [chitale2023task], allows for selection of higher rank during training without any overhead during inference. Task arithmetic assumes that fine-tuning a pre-trained model pushes the weights away towards a semantically relevant direction in the weight space, thus enabling model compression using simple arithmetic operations with task vectors [chitale2023task]. However, task arithmetic suffers from number of shortcomings including finding optimal scaling coefficients and suboptimal performance in the absence of a memory buffer [zhou2024metagpt].
@@
Typically, when using arbitrary weights, a higher rank usually facilitates learning new tasks but tends to increase forgetting with significant increase in number of parameters, while lower-rank mitigates forgetting with limited adaptation [lu2024adaptive, biderman2024lora]. There often exists a balance between plasticity and stability at a moderately small (but not too small) rank, which may enhance the benefits of CL [lu2024adaptive]. To this end, we introduce PEARL (Parameter-Efficient Adaptive Ranking for lifelong learning with LoRA), a rehearsal-free, parameter-isolation-based PEFT vision CL framework that mitigates catastrophic forgetting through dynamic low-rank adaptation. Specifically, PEARL leverages the information contained in the reference task weights  [^1] We use reference task interchangeably with pre-training when available. Otherwise, it refers to the first task in CL. and dynamically determines the rank of task-specific, layer-specific LoRA components during training based on the current task’s proximity to reference task weights in parameter space. With our dynamic threshold criterion, a close proximity would result in a lower rank allocation and maximum information re-use while a distant proximity results in a larger rank selection, thereby finding an optimal balance between learning and forgetting. To demonstrate the versatility of PEARL, we perform extensive analyses across various datasets, multiple CL scenarios and different architectures, an overview of which can be found in Appendix A.1. Our results demonstrate that PEARL variants consistently outperform the evaluated baselines in all scenarios by a huge margin. Our contributions are as follows:
@@ 2 Related Works
Exemplar and Non-Exemplar Based CL: Experience Rehearsal (ER), which stores and replays samples from previous tasks, has remained a leading approach to addressing the issue of catastrophic forgetting in CL. Several works build upon ER: DER++ [buzzega2020dark] proposed a consistency regularization technique to enhance stability and generalization. Guided by the GWT, TAMiL [bhat2022task] and AGILE [bhat2024mitigating] incorporate information reuse and task-specific, parameter-efficient adapters that re-project the backbone features onto a global workspace. However, these approaches face additional computational and memory overhead due to model surrogates and experience rehearsal. Since storing samples from previous tasks is not always feasible, non-exemplar-based approaches are gaining traction. Non-Exemplar Class-IL (NECIL) is particularly important in scenarios where data confidentiality is crucial due to privacy or security concerns, and where data storage has a limited lifespan. Early contributions to this domain include LwF [li2017learning] and EWC [kirkpatrick2017overcoming], both of which employ regularization techniques to reduce catastrophic forgetting. More recent methods, such as PASS [zhu2021prototype] and IL2A [zhu2021class], have further advanced NECIL by generating prototypes for previous classes without the need to retain the original images. SSRE [zhu2022self] introduced a reparameterization method to balance old and new knowledge, along with additional self-training that utilizes external data as an alternative to exemplars. FeTrIL [petit2023fetril] proposed a framework that combines a fixed feature extractor with a pseudo-feature generator using geometric transformations to achieve a better stability-plasticity balance. FeCAM [goswami2024fecam] addressed the limitations of Euclidean distance when feature distributions are non-stationary and suggested using anisotropic Mahalanobis distance as a substitute for Euclidean distance when employing a frozen feature extractor for prototype classification. Overall, the NECIL benchmark presents a significant challenge as it prohibits the storage of any samples from previous tasks while requiring effective performance in a Class-IL setting. PEARL also aligns with the NECIL benchmark as it operates without any experience rehearsal.
@@
Parameter Efficient Fine-Tuning (PEFT) techniques that have emerged in the space of Large Language Models have been quite pivotal in optimizing model performance with little memory and computational overhead. Seminal works such as Adapter modules [houlsby2019parameter], Prompt Tuning [brown2020language], BitFit [zaken2021bitfit], LoRA [hu2021lora], and DoRA [liu2024dora] have demonstrated the efficacy of selective fine-tuning all the while preserving model generalization and enhancing model adaptability. Specifically, LoRA introduces two compact, trainable low rank matrices into each layer’s weight matrix, allowing model to adapt to new tasks with minimal memory and computational footprint. Within CL, several approaches leverage PEFT techniques to mitigate catastrophic forgetting: Prompt-based methods such as L2P [wang2022learning] and DualPrompt [wang2022dualprompt] introduce task-specific prompts to facilitate learning new tasks while preserving consolidated knowledge. S-Prompt [wang2022s] and CODA-Prompt [smith2023coda] extend these works by employing structural prompts to map discriminative relationships and through Schmidt orthogonalization respectively. In order to maintain an optimal balance between computational complexity and accuracy, adapter-based methods such as SEMA [SEMA] and DIA [DIA] propose self-expanding, dynamic integration of task-specific adapters. Similarly, EASE [zhou2024expandable] creates task-specific sub-spaces by employing lightweight adapter modules for each new task, thereby maintaining resource efficiency. In contrast to adapter-based methods, InfLoRA [liang2024inflora] inserts a small number of LoRA parameters to re-parameterize the pre-trained weights and shows that fine-tuning these inserted parameters is equivalent to fine-tuning the pre-trained weights within a subspace. C-LoRA [DBLP:journals/corr/abs-2304-06027] incorporates a self-regularization strategy, which penalizes updates to parameters already modified by previous tasks. This mechanism helps preserve prior knowledge and enhances stability across sequential tasks. While these approaches have shown relative success, their implementation is tailored specifically to pre-trained vision transformers and do not consider task dynamics, limiting the potential for information reuse and dynamic resource allocation.
@@
Task Arithmetic [ilharcoediting], centered around task vectors, offers a new paradigm for steering the behavior of pre-trained / reference network towards multiple downstream tasks using task similarity. Inspired by recent works on weight interpolation [frankle2020linear, ilharco2022patching, ainsworth2022git], task vectors are defined to be the element-wise difference between pre-trained / reference network weights before and after fine-tuning on a particular task. A task vector specifies a direction in the weight space of a pre-trained model, such that movement in that direction improves performance on a given task [ilharcoediting]. As these task vectors can be modified and combined together in simple arithmetic operations, new tasks can be learnt via addition while interfering tasks can be forgotten via negation. Inspired by these findings, chitale2023task proposed an approach to continually train transformer-based vision models using low-rank adaptation and task arithmetic. Although task arithmetic is quite lucrative for its simplicity and intuitiveness, it suffers from several shortcomings such as finding optimal scaling coefficients and sub-optimal performance in the absence of memory buffer to finetune the network after weight aggregation [zhou2024metagpt].
@@ 3.1 Information aggregation during reference task training
where b{b} is a training batch, $\mathcal{L}_{ce}$ is cross-entropy loss, $\mathcal{D}_{r}$ is the data distribution for the reference task, and $\sigma(.)$ is the softmax function. Since pre-trained models demonstrate strong generalization capabilities, we expect to aggregate generic information that can be re-used across tasks later during CL. In scenarios where pre-training on a large corpus of data is not a possibility, the first task in a CL setup can also be used as a reference task.
@@ 3.2 Information re-use and dynamic resource allocation during CL
where $\|\cdot\|$ be a unitarily invariant norm. The Eckart-Young-Mirsky Theorem (3.1) implies that the best low k-rank approximation of a matrix can be achieved by selecting top-k singular values and such an approximation maintains the least error in terms of unilaterally invariant norms including the Frobenius and spectral norms. Our novelty lies in finding the best-k that eliminates the need for extensive hyperparameter tuning of the desired k and provides an optimum trade-off between performance and growth in number of learnable parameters. To select the best-k, we propose a dynamic threshold on the cumulative explained variance of singular values as follows:
@@
where $\mathcal{T}$ represents a normalized squared error between current task and reference task weights, and $\circ$ represents element-wise multiplication. Through $\mathcal{T}$, we approximate the proximity of the current task with respect to the reference task in the parameter space. $\mathcal{T}$ is not guaranteed to lie in $[0,1]$ and $\mathcal{T}$ can be as large as $2$ in extreme cases. Therefore, whenever $\mathcal{T}>1$, we simply clip it to 1 in our implementation. Once $\mathcal{T}$ is computed, the dynamic rank for the task-specific, layer-specific LoRA can be found as follows:
$k=\arg\max_{j}\left(\frac{\sum_{i=1}^{j}\mathbf{\sigma}_{i}^{2}}{\sum_{i=1}^{r}\mathbf{\sigma}_{i}^{2}}\geq\mathcal{T}\right)+1$ (4)
Equation 3 and 4 ensure that the k is bounded and select the rank in proportion to the cumulative explained variance of singular values. The +1+1 term in Equation 4 is to ensure the chosen rank is at least one singular value, preventing a degenerate zero-rank projection and preserving minimal expressive capacity. Subsequently, the low-rank matrices can be initialized from the SV-Decomposition of $\mathcal{W}^{t}_{c}$ (flattened and re-shaped to 2D if necessary) as follows:
@@ 3.3 Putting it all together
Table 1: Comparison of Class-IL performance of exemplar-based and NECIL approaches across different datasets and number of tasks. Baselines use randomly initialized ResNet-18 as a backbone. Exemplar-based approaches use a buffer size of 200. NECIL results are taken from [goswami2024resurrecting], while for the rest, we report an average of 3 runs. Best results are marked in bold while second best are underlined.
@@
DER++ [buzzega2020dark] NeurIPS 2020 28.8 47.89 25.42 43.94 17.75 35.89 10.15 27.26
TAMiL [bhat2022task] ICLR 2023 41.43 58.09 32.20 49.87 26.32 43.91 20.46 40.94
AGILE [bhat2024mitigating] CoLLAs 2024 45.73 57.66 33.23 42.88 20.14 43.79 20.19 43.81
LwF [li2017learning] ECCV 2016 45.35 61.94 26.14 46.14 38.81 49.70 27.42 38.77
IL2A [zhu2021class] NeurIPS 2021 - - 31.70 48.40 - - 25.30 42.01
PASS [zhu2021prototype] CVPR 2021 49.75 63.39 37.78 52.18 36.44 48.64 26.58 38.65
SSRE [zhu2022self] CVPR 2022 42.39 56.57 29.44 44.38 30.13 43.20 22.48 34.93
FeTrIL [petit2023fetril] WACV 2023 45.11 60.42 36.69 52.11 29.91 43.99 23.88 36.35
FeCAM [goswami2024fecam] NeurIPS 2024 47.28 61.37 33.82 48.58 25.62 39.85 23.21 35.32
@@ 4 Results
Experimental settings: We evaluate PEARL across two distinct scenarios of CL: Class-IL and Task-IL, a complete overview of which can be found in Appendix A.1. To facilitate this evaluation, we consider a comprehensive array of baseline methods: (i) exemplar-based approaches that involve information reuse and task-specific adapters (e.g., TAMiL, AGILE), (ii) NECIL approaches (e.g., PASS, FeTriL, FeCAM), (iii) parameter-isolation approaches that operate within a fixed capacity (e.g., NISPA) and those that expand beyond it (e.g., PNN, PAE), and (iv) PEFT approaches for CL (e.g., L2P, C-LoRA, InfLoRA, SEMA, ADA). We also provide a justification for the choice of these baseline in Appendix A.2. The first three categories utilize randomly initialized ResNet-18 as their backbone, whereas PEFT approaches leverage the ViT-Base-16 model pre-trained on ImageNet-21k. It is important to note that pre-training is often a luxury for smaller, custom models. Consequently, in accordance with the training protocols of the baseline methods, we employ randomly initialized convolutional models, while the ViT model is pre-trained. To demonstrate the versatility of PEARL across different architectures, we explore three distinct backbones: (i) Convolutional networks such as ResNet-18 (R18) and a ResNet-18-like architecture incorporating Blueprint Separable Convolutions (BSC) [li2022blueprint], and (ii) the Vision Transformer (ViT) Vit-Base-16 pre-trained on ImageNet-21k. For each subsequent task, we adapt convolutional layer, point-wise convolutional layer and linear layer projecting key in self-attention using dynamic low-rank adaptation in each of these backbones respectively. More information on backbones and their settings can be found in Appendix C.
@@
Comparison with convolutional CL approaches: Table 1 presents a comparison of PEARL against exemplar-based and NECIL approaches across four different Class-IL settings. We report results for both of our convolutional variants. Since the backbone is randomly initialized without pre-training, PEARL convolutional variants use the first CL task as the reference task and assign LoRA components for subsequent tasks based on their proximity to the first task in the parameter space (Algorithm 2). As shown, PEARL (R18) outperforms all considered baselines without relying on experience rehearsal while exhibiting modest model growth with an increasing number of tasks (see Table 11). Although a larger buffer size might enhance the performance of rehearsal-based approaches, any improvements would likely be outweighed by the increased computational and memory costs [TriRE2023]. Additionally, NECIL approaches do not always re-use information across tasks, which often leads to subpar performance. In contrast, PEARL capitalizes on the information contained in the first task for every subsequent task, achieving an optimal balance between performance and model size. While the absence of a pre-trained model may limit the generalizability of the first task, the framework’s ability to reuse information and dynamically allocate resources allows PEARL to efficiently assimilate information from both the first and current tasks, resulting in superior performance across tasks.
@@
Table 2: Class-IL performance comparison on ImageNet-R across 5, 10, and 20 task sequences. Baseline results with standard deviation are taken from [liang2024inflora] and that of PEARL are averaged over 3 runs with errors indicating standard deviation across these runs. The rest of the results are taken from their respective manuscripts. Best results are highlighted in bold. SeqLoRA is a sequential learning baseline which fine-tunes LoRA components on multiple tasks without any additional support.
@@
L2P [wang2022learning] CVPR 2022 64.13 $\pm$ 0.78 68.66 $\pm$ 0.41 62.54 $\pm$ 0.24 67.98 $\pm$ 0.27 57.92 $\pm$ 0.28 64.57 $\pm$ 0.29
DualPrompt [DBLP:conf/eccv/0002ZESZLRSPDP22] ECCV 2022 67.88 $\pm$ 0.17 71.16 $\pm$ 0.31 65.41 $\pm$ 0.52 69.39 $\pm$ 0.43 61.00 $\pm$ 0.72 65.80 $\pm$ 0.67
CODA-P [smith2023coda] CVPR 2023 73.09 $\pm$ 0.21 76.91 $\pm$ 0.21 71.47 $\pm$ 0.35 75.82 $\pm$ 0.29 67.28 $\pm$ 0.30 72.34 $\pm$ 0.17
C-LoRA [DBLP:journals/corr/abs-2304-06027] CoRR 2023 75.85 $\pm$ 0.31 78.85 $\pm$ 0.34 71.89 $\pm$ 0.45 75.33 $\pm$ 0.28 65.71 $\pm$ 0.60 70.63 $\pm$ 0.85
LAE [gao2023unified] ICCV 2023 73.84 $\pm$ 0.14 77.29 $\pm$ 0.45 71.70 $\pm$ 0.39 76.71 $\pm$ 0.10 66.98 $\pm$ 0.35 73.72 $\pm$ 0.05
InfLoRA [liang2024inflora] CVPR 2024 77.52 $\pm$ 0.37 82.01 $\pm$ 0.12 75.65 $\pm$ 0.14 80.82 $\pm$ 0.24 71.01 $\pm$ 0.45 77.28 $\pm$ 0.45
EASE [zhou2024expandable] CVPR 2024 70.80     - 73.37     - 78.80     -
SEMA [SEMA] CVPR 2025 79.78 84.75 78.00 83.56 74.53 81.75
DIA [DIA] CVPR 2025     -     - 79.03 85.61 76.32 83.51
TSVD [TSVD] ICLR 2025 73.58     - 74.55     - 79.13     -